# Single-file dynamics with general charge measures

### 25 Jan 2024 v1

@@ Abstract
We study charge fluctuations in single-file dynamics with general charge measures. The exact finite-time distribution of charge fluctuations is obtained in terms of a dressing transformation acting on the finite-time distribution of particle fluctuations. The transformation is mapped to a simple substitution rule for corresponding full-counting statistics. By taking the asymptotics of the dressing transformation we analyze typical and large scale charge fluctuations. Typical charge fluctuations in equilibrium states with vanishing mean charge are anomalous while large charge fluctuations undergo first and second order dynamical phase transitions out of equilibrium.
@@ 1 Introduction
A natural extension to the study of typical fluctuations is given by large deviation theory, characterizing large fluctuations, of order $\mathcal{O}(n)$, away from the mean. A basic problem of large deviation theory is to establish a large deviation principle for tail probabilities, i.e. showing that large fluctuations are suppressed exponentially in $n$ and to characterize the rate of suppression, encapsulated in the so-called rate function. The central limit theorem can then often be understood as the quadratic expansion of the rate function around the mean value.
The theory of large deviations is the natural language to describe the central notion of concentration of measure in statistical mechanics, see [2] for a review. Fluctuations of conserved quantities in many-body systems are captured by the so-called full-counting statistic of corresponding time-integrated current densities, a dynamical analogue of the partition function in equilibrium statistical mechanics. A time-integrated current is the many-body analogue of the sum of a sequence of variables, with the proviso that the current densities at different times are now correlated by the dynamics with randomness necessarily entering even for deterministic dynamics due to sampling from an ensemble of initial conditions.
This naturally leads to the problem of characterizing current fluctuations on both the large and typical scale. The question of large fluctuations in ergodic systems with a single conserved quantity has been thoroughly addressed by the development of macroscopic fluctuations theory, see [3] for a review. The theory establishes a large deviation principle and formulates a variational problem, the solution of which recovers the rate function. Recently, building upon the theory of generalized hydrodynamics [4, 5], an analogous ballistic macroscopic fluctuations theory [6] has been developed to describe large fluctuations in integrable systems. Similar results have also been derived by mapping the full-counting statistics in a non-equilibrium bipartite ensemble to a dual equilibrium problem using space-time duality [7].
@@
In Section 2, we introduce the notion of single-file dynamics and define the family of initial bipartite ensembles studied in the work. Attaching charge degrees of freedom to particles, we introduce inert charged single-file dynamics and comment on the role of charge inertness. We end the introductory section by giving a brief overview of the basic tools used in the study of fluctuations that we use in the remainder of the work.
@@ 2.1 Particle dynamics
$\varrho({\mathbf{x}},t_{0})=\lim_{\stackrel{{\scriptstyle Q_{p},L\to\infty}}{{Q_{p}/L=\textrm{const}}}}\varrho_{L}({\mathbf{x}},t_{0}).$ (4)
@@ 2.2 Inert charged dynamics
We attach a scalar charge $c\in\mathcal{C}\subseteq\mathbb{R}$ to each particle, sampled from a normalized measure, $\omega_{\mu}:\mathcal{C}\to\mathbb{R}^{+}_{0}$, $\int_{\mathcal{C}}{\rm d}\omega_{\mu}=1$, with a finite mean $|\int_{\mathcal{C}}{\rm d}\omega_{\mu}\,c=\mu|<\infty$ and variance $\int_{\mathcal{C}}{\rm d}\omega_{\mu}(c-\mu)^{2}=\sigma^{2}_{\omega}<\infty$. We consider measures whose support is not sign-definite, ${\rm s}_{\pm}\equiv{\rm supp}(\omega_{\mu})\cap\mathbb{R^{\pm}}\neq\emptyset$ and denote ±$\pm c_{\pm}\equiv\displaystyle\max_{c\in{\rm s}_{\pm}}\pm c$.
@@
$\varrho^{c}({\mathbf{x}},{\mathbf{c}},t_{0})=\lim_{\stackrel{{\scriptstyle Q_{p},L\to\infty}}{{Q_{p}/L=\textrm{const}}}}\varrho^{c}_{L}({\mathbf{x}},{\mathbf{c}},t_{0}).$ (7)
@@ 2.3 Inert charged single-file dynamics
$x_{1}(t)<x_{2}(t)<\ldots<x_{N-1}(t)<x_{N}(t),\quad\forall t\in\mathbb{R},$ (8)
which, due to charge inertness (5), is the same as imposing the single-file constraint on the particle dynamics $\phi_{s}$. We note that in an arbitrary crossing (i.e. one violating the single-file condition (8)) particle dynamics it is always possible to enforce the single-file constraint by adding an infinitely strong repulsive hardcore point interaction potential to the particles. For momentum conserving particle dynamics this leads to elastic particle scattering and reduces to permuting particle labels of colliding particles.
@@ 2.4 Full-counting statistics
The charge full-counting statistics (13) involves averaging over the composite initial measure (7) as charges are bound to particles. However, as inert charge do not effect particle dynamics, the inert particle full-counting statistics can be restricted to averaging solely over the particle initial measure (4)
@@ 3.1 Dressing the full-counting statistics
Consider the discrete $\mathbb{Z}_{2}$ charge measure
@@ 4 Dressing asymptotics
    The particle full-counting statistics obeys a large deviation principle with a strictly convex twice differentiable particle scaled cumulant generating function $F_{p}\in C^{2}(\mathbb{R})$
    $F_{p}(\lambda_{p})=\lim_{t\to\infty}t^{-\alpha}\log G_{p}(\lambda_{p}|t),$ (29)
@@
where $I_{p}(j_{p})\equiv-\lim_{t\to\infty}t^{-\alpha}\log\mathcal{P}_{p}(J_{p}=j_{p}t^{\alpha}|t)$ is the particle rate function. Since $F_{p}$ is assumed to be twice differentiable, the rate function $I_{p}$ is recovered by the Legendre-Fenchel transform [2]
@@
and is a twice differentiable strictly convex function on its domain $\mathcal{J}_{p}\equiv(j_{p}^{\rm min},j_{p}^{\rm max})$. We note that finiteness of scaled cumulants in (II) generally does not follow directly from the existence of the limit in (I), see [12] for a many-body system example. Instead, a sufficient condition is that the logarithm of the particle full-counting statistics $\log G_{p}(\lambda_{p}|t)$ is complex-analytic at the origin $\lambda_{p}=0$ for all $t$, see [8]. Assumption (II) should be understood as a regularity condition on the particle dynamics that ensures central limit behavior on the typical scale.
@@ 4.1 Typical charge fluctuations
Eq. (39) shows that the charge variance decomposes into a contribution stemming from typical particle fluctuations with variance $\sigma_{p}^{2}$ and average charge $\mu_{\overline{\nu}}$ and a contribution from the variance of the charge measure $|\sigma_{\omega}^{2}|$ carried by the average particle current $|s_{1}^{(p)}|$.
@@ 4.2 Large charge fluctuations
Typical charge fluctuations depend only on the first two moments of the charge measure, whereas large fluctuations are sensitive to the full global structure of the measure. In the following analysis it will be convenient that the Fourier transform of the measure $\hat{\omega}_{\mu}(k)$ is an entire function. By the Paley-Wiener theorem [26], a sufficient assumption is that the charge measure decays faster than exponentially on the real line
$|\omega_{\mu}(c)|<ae^{-b|c|}\quad{\rm as}\enspace|c|\to\infty,\quad\textrm{forfixed $a$ and all }b\in\mathbb{R^{+}}.$ (43)
@@
as a consequence of the current bounds (46-49). The Legendre-Fenchel transform (31) relates the derivative of the particle rate functions to the inverse of the derivative of the scaled cumulant generating function, $\partial_{j_{p}}I_{p}(j_{p})=(\partial_{\lambda_{p}}F_{p})^{-1}(j).$ Differentiability and strict convexity of $F_{p}$ on $\mathbb{R}$ then combine to give divergent derivatives of the particle rate function at the boundaries of $\mathcal{J}_{p}$
@@ 4.2.1 First optimization problem
with ${\rm i}\kappa_{\pm}\in{\rm i}\mathbb{R}$ the unique solutions of saddle point equations on the imaginary axis
@@
Having established these general properties of the split rate functions, we are in a position to describe the structure of the split charge rate functions $I_{c}^{\pm}$ resulting from the minimization (58). Note that the properties (61-64) are similar to those used in the analysis of the minimization problem (58) for a $\mathbb{Z}_{2}$ charge measure (27) in [20], which serves as a useful example and visual guide that contains most of the structure of the general problem studied here.
@@
Since $I^{\pm}_{c,p}$ are continuous strictly convex functions of $j_{p}$ whose derivatives change sign on $\mathcal{J}_{p}^{\mp}$, each function has a single minimum at $j_{p}=j^{\pm}_{\rm crit}(j_{c})$ in the interior of the domains $\mathcal{J}_{p}^{\mp}$ where the derivetive of $I^{\pm}_{c,p}$ vanish
@@ 4.2.2 Second optimization problem
The simplest case is that one of the split charge rate functions $I^{\pm}_{c}$ dominates the other on the entire charge current domain, trivially resulting in a differentiable charge rate function $I_{c}(j)=I_{c}^{\pm}(j_{c})$. On the other hand, nothing prevents one of the rate functions overtaking the other at a generic (i.e. non-zero) value of the charge current $j_{c}^{\rm cusp}\neq 0$
$I^{+}_{c}(j_{c}^{\rm cusp}\neq 0)=I^{-}_{c}(j_{c}^{\rm cusp}\neq 0),$ (72)
@@
We accordingly distinguish two cases. In the first case, the same function $I_{c}^{\pm}$ is smaller on both sides of the origin and the resulting $I_{c}$ is trivially non-differentiable at the origin and strictly convex in a neighborhood of the origin. In the second case, $I_{c}$ switches from $I^{\pm}_{c}$ to $I_{c}^{\mp}$ as it crosses the origin. The resulting $I_{c}$ is again non-differentiable at the origin while its strict convexity follows from the strict convexity of $I^{\pm}_{c,p}(j_{c}=0,j_{p})$ on $\mathcal{J}_{p}$ and (64).
To summarize, the charge rate functions $I_{c}$, obtained as solutions of the optimization problem (71), can exhibit two basic types of non-differentiable behavior. The first type correspond to a non-convex cusp in the rate function that occurs away from the origin, at which the two split rate functions exchange dominance. The second type is manifested as non-differentiability of the rate function at the origin. The split rate functions can either exchange dominance at the origin or on of the rate function dominates on both sides of the origin. Both cases preserve strict convexity of the charge rate function. We emphasize that the charge rate function can combine both a non-differentiable cusp (72) with non-differentiability at the origin (74).
@@ 4.3 Dynamical phase transitions
see also the discussion in [20] for a more elaborate discussion. It is well-known that scaled cumulant generating functions are convex and are as such recoverable from a rate function by the Legendre-Fenchel transform irrespective of its differentiability [2]
$F_{c}(\lambda_{c})=\inf_{j_{c}}\{\lambda_{c}j_{c}-I_{c}(j_{c})\}.$ (76)
A non-convex cusp (72) at $j_{c}^{\rm cusp}\neq 0$ of the charge rate function $I_{c}$ is accordingly mapped to a convex non-differentiable corner of $F_{c}$. From the miss-match of first derivatives of $F_{c}$ at the corner we conclude that the non-convex cusp of $I_{c}$ corresponds to a first-order dynamical phase transition.
On the other hand, a non-differentiable point at the origin of a strictly convex charge rate function $I_{c}$ (74) is mapped to a flat interval of $F_{c}$, whose boundaries are given by the right and left derivatives of the rate function at the origin
@@
where positivity follows from strict convexity of $I_{c}$. The mismatch of second derivatives shows that non-differentiability at the origin of the charge rate function (74) corresponds to a second order dynamical phase transition.

### 01 Feb 2024 v2

@@ Abstract
We study charge fluctuations in single-file dynamics with general charge measures. The exact finite-time distribution of charge fluctuations is obtained in terms of a dressing transformation acting on the finite-time distribution of particle fluctuations. The transformation is mapped to a simple substitution rule for corresponding full-counting statistics. By taking the asymptotics of the dressing transformation, we analyze typical and large scale charge fluctuations. Typical charge fluctuations in equilibrium states with vanishing mean charge are anomalous, while large charge fluctuations undergo first and second order dynamical phase transitions out of equilibrium.
@@ 1 Introduction
A natural extension to the study of typical fluctuations is given by large deviation theory, characterizing large fluctuations, of order $\mathcal{O}(n)$, away from the mean. A basic problem of large deviation theory is to establish a large deviation principle for tail probabilities, i.e. to show that large fluctuations are suppressed exponentially in $n$ and to characterize the rate of suppression, encapsulated in the so-called rate function. The central limit theorem can then often be understood as the quadratic expansion of the rate function around the mean value.
The theory of large deviations is the natural language to describe the central notion of concentration of measure in statistical mechanics both in and out of equilibrium [2, 3]. Fluctuations of conserved quantities in many-body systems are captured by the so-called full-counting statistic of corresponding time-integrated current densities, a dynamical analogue of the partition function in equilibrium statistical mechanics. A time-integrated current is the many-body analogue of the sum of a sequence of variables, with the proviso that the current densities at different times are now correlated by the dynamics with randomness necessarily entering even for deterministic dynamics due to sampling from an ensemble of initial conditions.
This naturally leads to the problem of characterizing current fluctuations on both the large and typical scale. The question of large fluctuations in ergodic systems with a single conserved quantity has been thoroughly addressed by the development of macroscopic fluctuations theory [4]. The theory establishes a large deviation principle and formulates a variational problem, the solution of which recovers the rate function. Recently, building upon the theory of generalized hydrodynamics [5, 6], an analogous ballistic macroscopic fluctuations theory [7] has been developed to describe large fluctuations in integrable systems. Similar results have also been derived by mapping the full-counting statistics in a non-equilibrium bipartite ensemble to a dual equilibrium problem using space-time duality [8].
@@
In Section 2, we introduce the notion of single-file dynamics and define the family of initial bipartite ensembles studied in the work. Attaching charge degrees of freedom to particles, we introduce inert charged single-file dynamics and comment on the role of charge inertness. We end the introductory section by giving a brief overview of the basic tools that we use in the remainder of the work.
@@ 2.1 Particle dynamics
$\varrho({\mathbf{x}},t_{0})=\lim_{\stackrel{{\scriptstyle Q_{p},L\to\infty}}{{Q_{p}/L=(\rho_{-}+\rho_{+})/2}}}\varrho_{L}({\mathbf{x}},t_{0}).$ (4)
@@ 2.2 Inert charged dynamics
We attach a scalar charge $c\in\mathcal{C}\subseteq\mathbb{R}$ to each particle, sampled from a normalized measure, $\omega_{\mu}:\mathcal{C}\to\mathbb{R}^{+}_{0}$, $\int_{\mathcal{C}}{\rm d}\omega_{\mu}=1$, with a finite mean $|\int_{\mathcal{C}}{\rm d}\omega_{\mu}\,c|=|\mu|<\infty$ and variance $\int_{\mathcal{C}}{\rm d}\omega_{\mu}(c-\mu)^{2}=\sigma^{2}_{\omega}<\infty$. We consider measures whose support is not sign-definite, ${\rm s}_{\pm}\equiv{\rm supp}(\omega_{\mu})\cap\mathbb{R^{\pm}}\neq\emptyset$ and denote ±$\pm c_{\pm}\equiv\displaystyle\max_{c\in{\rm s}_{\pm}}\pm c$.
@@
$\varrho^{c}({\mathbf{x}},{\mathbf{c}},t_{0})=\lim_{\stackrel{{\scriptstyle Q_{p},L\to\infty}}{{Q_{p}/L=(\rho_{-}+\rho_{+})/2}}}\varrho^{c}_{L}({\mathbf{x}},{\mathbf{c}},t_{0}).$ (7)
@@ 2.3 Inert charged single-file dynamics
$x_{1}(t)<x_{2}(t)<\ldots<x_{N-1}(t)<x_{N}(t),\quad\forall t,$ (8)
which, due to charge inertness (5), is the same as imposing the single-file constraint on the particle dynamics $\phi_{s}$. We note that in an arbitrary crossing (i.e. one violating the single-file condition (8)) particle dynamics it is always possible to enforce the single-file constraint by adding an infinitely strong repulsive hardcore point interaction to the particles. For momentum conserving particle dynamics this leads to elastic particle scattering and reduces to permuting particle labels of colliding particles.
@@ 2.4 Full-counting statistics
The charge full-counting statistics (13) involves averaging over the composite initial measure (7) as charges are bound to particles. However, since inert charge do not effect particle dynamics, the inert particle full-counting statistics can be restricted to averaging solely over the particle initial measure (4)
@@ 3.1 Dressing the full-counting statistics
We note that $\hat{\omega}_{\mu_{\mp}}(\pm{\rm i}\lambda_{c})=\int_{\mathcal{C}}{\rm d}\omega_{\mu_{\mp}}e^{\pm c\lambda_{c}}$ are the charge full-counting statistics of a single particle going left-to-right or right-to-left respectively. Since charges are independent in the initial measure (6), the corresponding $n$-particle result factorizes in terms of the single-particle one. The substitution rule (26) therefore dresses the particle full-counting statistics by adjoining the left/right-going charge-full counting statistics with fixed particle number to the corresponding particle crossings.
@@
Consider the $\mathbb{Z}_{2}$ charge measure
@@ 4 Dressing asymptotics
    The particle full-counting statistics satisfy a large deviation principle with a strictly convex twice differentiable particle scaled cumulant generating function $F_{p}\in C^{2}(\mathbb{R})$
    $F_{p}(\lambda_{p})\equiv\lim_{t\to\infty}t^{-\alpha}\log G_{p}(\lambda_{p}|t),$ (29)
@@
where $I_{p}(j_{p})\equiv-\lim_{t\to\infty}t^{-\alpha}\log\mathcal{P}_{p}(J_{p}=j_{p}t^{\alpha}|t)$ is the particle rate function and $\asymp$ indicates asymptotic logarithmic equivalence. Since $F_{p}$ is assumed to be twice differentiable, the rate function $I_{p}$ is recovered by the Legendre-Fenchel transform [3]
@@
and is a twice differentiable strictly convex function on its domain $\mathcal{J}_{p}\equiv(j_{p}^{\rm min},j_{p}^{\rm max})$. We note that finiteness of scaled cumulants in (II) generally does not follow directly from the existence of the limit in (I), see [13] for a many-body system example. Instead, a sufficient condition is that the logarithm of the particle full-counting statistics $\log G_{p}(\lambda_{p}|t)$ is complex-analytic at the origin $\lambda_{p}=0$ for all $t$, see [9]. The assumption (II) should be understood as a regularity condition on the particle dynamics that ensures central limit behavior on the typical scale.
@@ 4.1 Typical charge fluctuations
Eq. (39) shows that the charge variance decomposes into a contribution stemming from typical particle fluctuations with variance $\sigma_{p}^{2}$ and average charge $\mu_{\overline{\nu}}$ and a contribution from the variance of the charge measure $\sigma_{\omega}^{2}$ carried by the average particle current $|s_{1}^{(p)}|$.
@@ 4.2 Large charge fluctuations
Typical charge fluctuations depend only on the first two moments of the charge measure, whereas large fluctuations are sensitive to the full structure of the measure. In the following analysis it will be convenient that the Fourier transform of the measure $\hat{\omega}_{\mu}(k)$ is an entire function. By the Paley-Wiener theorem [27], a sufficient assumption is that the charge measure decays faster than exponentially on the real line
$\omega_{\mu}(c)<ae^{-b|c|}\quad{\rm as}\enspace|c|\to\infty,\quad\textrm{for fixed $a$ and all }b\in\mathbb{R^{+}}.$ (43)
@@
as a consequence of the current bounds (46-49). The Legendre-Fenchel transform (31) relates the derivative of the particle rate functions to inverse derivatives of the scaled cumulant generating function, $\partial_{j_{p}}I_{p}(j_{p})=(\partial_{\lambda_{p}}F_{p})^{-1}(j).$ Differentiability and strict convexity of $F_{p}$ on $\mathbb{R}$ then combine to give divergent derivatives of the particle rate function at the boundaries of $\mathcal{J}_{p}$
@@ 4.2.1 First optimization problem
with $\kappa_{\pm}\in\mathbb{R}$ the unique solutions of saddle point equations on the imaginary axis
@@
Having established these general properties of the split rate functions, we are in a position to describe the structure of the split charge rate functions $I_{c}^{\pm}$ resulting from the minimization (58). We note that the properties (61-64) are similar to those used in the analysis of the minimization problem (58) for a $\mathbb{Z}_{2}$ charge measure (27) in [21], which serves as a useful example and visual guide that contains most of the structure of the general problem studied here.
@@
Since $I^{\pm}_{c,p}$ are continuous strictly convex functions of $j_{p}$ whose derivatives change sign on $\mathcal{J}_{p}^{\mp}$, each function has a single minimum at $j_{p}=j^{\pm}_{\rm crit}(j_{c})$ in the interior of the domains $\mathcal{J}_{p}^{\mp}$ where the derivatives of $I^{\pm}_{c,p}$ vanish
@@ 4.2.2 Second optimization problem
The simplest case is that one of the split charge rate functions $I^{\pm}_{c}$ dominates the other on the entire charge current domain, trivially resulting in a differentiable charge rate function $I_{c}(j)=I_{c}^{\pm}(j_{c})$. On the other hand, nothing prevents one of the rate functions overtaking the other at a generic value of the charge current $j_{c}^{\rm cusp}$
$I^{+}_{c}(j_{c}^{\rm cusp})=I^{-}_{c}(j_{c}^{\rm cusp}),$ (72)
@@
We accordingly distinguish two cases. In the first case, the same function $I_{c}^{\pm}$ is smaller on both sides of the origin and the resulting $I_{c}$ is trivially non-differentiable at the origin and strictly convex in a neighborhood of the origin. In the second case, $I_{c}$ switches from $I^{\pm}_{c}$ to $I_{c}^{\mp}$ as it crosses the origin. The resulting $I_{c}$ is again non-differentiable at the origin but is no longer necessarily convex and can exhibit a non-convex cusp.
To summarize, the charge rate functions $I_{c}$, obtained as solutions of the optimization problem (71), can exhibit two basic types of non-differentiable behavior. The first type correspond to a non-convex cusp in the rate function, at which the two split rate functions exchange dominance. The second type is manifested as non-differentiability of the rate function at the origin. In case one of the split rate functions dominates on both sides of the origin the charge rate function is strictly convex. When split rate functions exchange dominance at the origin the charge rate function can exhibit a non-convex cusp at the origin. We emphasize that the charge rate function can combine both a non-differentiable cusp (72) for $j_{c}^{\rm cusp}\neq 0$ with non-differentiability at the origin (74).
@@ 4.3 Dynamical phase transitions
commonly referred to as the dynamical free energy [2], see also the discussion [21].
Figure 1: A first order dynamical phase transition. (a) Split charge rate functions $I_{c}^{\pm}$ (red/blue) entering the second minimization problem (71). The charge rate function $I_{c}$ (solid line) has non-convex cusp at $j_{c}^{\rm cusp}$ (gray circle). (b) Split charge scaled cumulant generating functions $F_{c}^{\pm}(\lambda_{c})$ (77) (red/blue) entering the second optimization problem (78). The non-convex cusp of $I_{c}$ is mapped to a non-differentiable corner (gray circle) of the charge scaled cumulant generating function $F_{c}$ (solid line) at $\lambda_{c}^{\rm cor}$.
Scaled cumulant generating functions are convex and are as such recoverable from a rate function by the Legendre-Fenchel transform irrespective of its differentiability [3]
$F_{c}(\lambda_{c})=\sup_{j_{c}}\{\lambda_{c}j_{c}-I_{c}(j_{c})\}.$ (76)
The two-step structure of the charge rate function is inherited by the optimization problem (76), leading to an initial optimization for the split charge scaled cumulant generating functions
$F^{\pm}_{c}(\lambda_{c})=\sup_{j_{c}}\{\lambda_{c}j_{c}-I_{c}^{\pm}(j_{c})\},$ (77)
followed by picking the larger of the two
$F_{c}(\lambda_{c})=\max\{F_{c}^{-}(\lambda_{c}),F_{c}^{+}(\lambda_{c})\}.$ (78)
A non-convex cusp (72) at $j_{c}^{\rm cusp}$ (including at $j_{c}^{\rm cusp}=0$) of the charge rate function $I_{c}$ is accordingly mapped to a convex non-differentiable corner of $F_{c}$ at $\lambda_{c}^{\rm cor}$, see Figure 1. From the miss-match of first derivatives of $F_{c}$ at the corner we conclude that a non-convex cusp of $I_{c}$ corresponds to a first-order dynamical phase transition.
Figure 2: A second order dynamical phase transition. (a) A pair of split charge rate functions $I_{c}^{\pm}$ (red/blue) exchange dominance at the origin $j_{c}=0$, where the strictly convex charge rate function $I_{c}$ (solid line) has non-differentiable point (gray circle). (b) Split charge scaled cumulant generating functions $F_{c}^{\pm}(\lambda_{c})$ (77) (red/blue) exhibit a flat plateau. The non-convex point of $I_{c}$ at the origin is mapped to a flat interval (79) (gray line) of the charge scaled cumulant generating function $F_{c}$ (solid line) with a discontinuous second derivative at the boundaries $\lambda_{c}^{\pm}$ (gray circles).
On the other hand, a non-differentiable point at the origin of a strictly convex charge rate function $I_{c}$ (74) is mapped to a flat interval of $F_{c}$, see Figure 2, whose boundaries are given by the right and left derivatives of the rate function at the origin
@@
where positivity follows from strict convexity of $I_{c}$. The mismatch of second derivatives shows that non-differentiability at the origin of a strictly convex charge rate function corresponds to a second order dynamical phase transition.